# Sonja Hohloch  &\&&  Pedro Santos

### v1

@@ 4.3. A semitoric system induced by the octagon
In De Meulenaere & Hohloch [de2021family, Section $3.1$] it is shown that $M=\tilde{L}^{-1}(0)/N$, where
(4.5) $\tilde{L}^{-1}(0):=\left\{(z_{1},...,z_{8})\in\mathbb{C}^{8}\left|\ \begin{aligned} &|z_{1}|^{2}+|z_{5}|^{2}=6\\ &|z_{2}|^{2}+|z_{5}|^{2}+|z_{7}|^{2}=10\\ &|z_{3}|^{2}+|z_{7}|^{2}=6\\ &|z_{4}|^{2}-|z_{5}|^{2}+|z_{7}|^{2}=4\\ &|z_{5}|^{2}-|z_{6}|^{2}+|z_{7}|^{2}=2\\ &|z_{5}|^{2}-|z_{7}|^{2}+|z_{8}|^{2}=4\end{aligned}\right.\right\}$
@@
Points on $(M,\omega)$ are written as equivalence classes of the form $[z]=[z_{1},...,z_{8}]$ with $z_{k}=x_{k}+iy_{k}\in\mathbb{C}$ for $1\leq k\leq 8$. The momentum map of the toric system $(M,\omega)$ is given as follows:
@@
(De Meulenaere &\&& Hohloch, [de2021family, Theorem 1.1]) Let $F=(J,H):M\rightarrow\mathbb{R}^{2}$ be given by
@@
(De Meulenaere &\&& Hohloch, [de2021family, Theorem 1.2]) Set $F_{t}:=(J,H_{t}):=(J,(1-2t)H+t\gamma X):(M,\omega)\rightarrow\mathbb{R}^{2}$. Then, $(M,\omega,F_{t})_{0\leq t\leq 1}$ is toric for $t=0$, of toric type for $0<t<t^{-},$ semitoric for $t^{-}<t<t^{+}$, and again of toric type for $t^{+}<t\leq 1$ where
$10<t^{-}:=\frac{1}{2(1+24\gamma)}<\frac{1}{2}<t^{+}:=\frac{1}{2(1-24\gamma)}<$ 1 .
@@
(De Meulenaere &\&& Hohloch, [de2021family, Proposition 3.7]) The focus-focus points of the semitoric systems $(M,\omega,F_{t})$, for $t^{-}<t<t^{+}$ are
@@
(De Meulenaere &\&& Hohloch, [de2021family, Proposition 1.3]) At $t=\frac{1}{2}$, the system $F_{\frac{1}{2}}$ has precisely two focus-focus fibers, each of which contains precisely two focus-focus points so that each of these two fibers has the shape of a double pinched torus.
@@
of $(\mathbb{C}^{8},\omega_{0})$. Recall that $(M,\omega)$ is obtained as symplectic reduction of $(\mathbb{C}^{8},\omega_{0})$ on a level set $\tilde{L}^{-1}(0)$, see Equation (4.5), by a torus action of $\mathbb{T}^{6}$, see Equation (4.6). Note that $\Psi_{1}$ and $\Psi_{2}$ preserve $\tilde{L}^{-1}(0)$ and are invariant under the action, hence they define symplectomorphisms on $(M,\omega)$. The result follows from considering the projections of $\Psi_{1},\Psi_{2}$,$\Psi_{1}\circ\Psi_{2}$ and $\Psi_{2}\circ\Psi_{1}$ to $(M,\omega)$. The symplectomorphisms do not depend on $t$. ∎
@@ 4.3.1. Symplectic topology of the fibers of the toric system
Application of Lemma 2.22: depending on the value one can use the probe with direction (±$(\pm 1,0)$ or $(0,\pm 1)$, see Figure 4.8. ∎
@@
We present the argument for the fiber over $(\frac{3}{2},\frac{3}{2})$, the other ones are analogous. See Borman [borman2013quasi] and Abreu & Macarini [abreu2013remarks] for more details on the general method.
  • Notice that $(M,\omega)$ can be viewed as the symplectic reduction of $\mathbb{CP}^{2}\times\mathbb{C}\mathbb{P}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}$ for an appropriate choice of subtorus with associated Lie algebra vectors and appropriate chosen symplectic volume of $\mathbb{CP}^{2}$ and $\mathbb{CP}^{1}$, see Figure 4.9. For more details see Section 2.10 and Abreu & Macarini [abreu2013remarks].
  • Notice that the central fiber in $(\mathbb{CP}^{1},\omega_{0})$ and the fiber over $(\frac{1}{3},\frac{1}{3})$ in $(\mathbb{CP}^{2},\omega_{FS})$ are stems, and in particular superheavy for any quasi-state, see Polterovich & Rosen [polterovich2014function, Proposition $6.1.13$].
  • For each $\mathbb{CP}^{1}$ and for $\mathbb{CP}^{2}$ we choose quasi-states as in Borman [borman2013quasi, Theorem $3.1$] to create a product quasi-state in $\mathbb{CP}^{2}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}$, see Borman [borman2013quasi, Corollary $3.2$], such that the product of the above mentioned fibers is superheavy with respect to this quasi state. Furthermore, this quasi-state satisfies the PB-inequality.
@@
Figure 4.9. $(M,\omega)$ as the symplectic reduction of $\mathbb{CP}^{2}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}$. The dotted lines come in pairs, and each correspond to a reduction by $\mathbb{CP}^{1}$. The triangle corresponds to $\mathbb{CP}^{2}$.
@@ 4.3.2. Symplectic topology of the fibers of the semitoric systems
Figure 4.10. Representative of the polytope invariant for the semitoric system when $\vec{\epsilon}=(-1,1,-1,1)$. The dotted lines represent the eigenrays associated with the focus-focus fibers. The red circles represent the $4$ different focus-focus values.
@@ 4.3.3. Case $10<h<$ 1
[…]
@@ 4.3.5. The case $h=1$
  • Applying Lemma 2.22 to Figure 4.16 2.22 we conclude that every fiber of the form $(h_{\vec{\epsilon}}\circ F_{t})^{-1}(1,b)$ with $21<b<$ 2, where $h_{\vec{\epsilon}}$ is the straightening homeomorphism associated with the polytope of Figure 4.16, is displaceable.
@@ Appendix A Polytope invariant for the semitoric system induced by the octagon
Let $(M,\omega)$ be a closed symplectic manifold of dimension $2n$. The Liouville measure on $(M,\omega)$ is the measure induced by the volume form $\frac{1}{(2\pi)^{n}}\frac{\omega^{n}}{n!}$. The Duistermaat-Heckman measure $\mu_{J}$ for a Hamiltonian $J:M\rightarrow\mathbb{R}$ inducing an $S^{1}$-action is defined as $\mu_{J}([a,b])=\text{vol}(J^{-1}([a,b]))$ where vol is taken with respect to the Liouville measure in $M$. By Duistermaat & Heckman [duistermaat1982variation]
@@
where the density $\rho_{J}(x)$, sometimes called the Duistermaat-Heckman function, is a continuous function, equal to the symplectic volume of the reduced orbifold $J^{-1}(x)/S^{1}$.
Let $J$ be the momentum map of an effective Hamiltonian $S^{1}$-action on $(M,\omega)$. For each subgroup $G\subset S^{1}$, let $M^{G}$ be the set of points in $M$ whose stabilizer is $G$. The connected components of $M^{S^{1}}$ are symplectic submanifolds, hence either points or surfaces, see Karshon [karshon1996periodic] for more details.
@@
(Chaperon, [chaperon1983quelques]) Let $(M,\omega,J)$ be as above. For each $p\in M^{S^{1}}$ there exist neighborhoods $U\subset M$ of $p$, $U_{0}\subset\mathbb{C}^{2}$ of $(0,0)$, and a symplectomorphism $\Psi:(U,\omega)\rightarrow(U_{0},\omega_{0})$, where $\omega_{0}=\frac{i}{2}(dz_{1}\wedge d\overline{z}_{1}+dz_{2}\wedge d\overline{z}_{2})$ making the following diagram commute
@@
(Vũ Ngǫc, [vu2007moment, Theorem 3.8]) Let $(M,\omega,\Phi)$ be a semitoric system and $f_{\vec{\epsilon}}$ a straightening homeomorphism associated with a representative of the polytope invariant. Furthermore, let $x\in\mathbb{R}$ be a regular value of $J$. If $\alpha^{+}(x)$ (resp. $\alpha^{-}(x)$) denotes the slope of the top (resp. bottom) boundary of the polytope $f_{\vec{\epsilon}}\circ\Phi(M)$, then the derivative of the Duistermaat-Heckman function is
@@
where $a^{\pm},b^{\pm}$ are the isotropy weights for the $S^{1}$-action at the corresponding vertices.
With Theorem A.2, we are able to compute representatives of the polytope invariant (see Section 2.8) for the semitoric system given in Section 4.3. We focus on the case of $4$ focus-focus values $(A,B,C,D)$. Furthermore, recall that there are $4$ elliptic-elliptic values. We label them $e_{1}$, $e_{2}$, $e_{3}$, $e_{4}$. We now compute a representative of the polytope invariant for the case $\vec{\epsilon}=(1,-1,1,-1)$, the other cases are analogous.
@@
Consider the image of $F(M)$ under the straightening homeomorphism $f_{\vec{\epsilon}}$. By abuse of notation we identify each $e_{i}$, $i=1,..,4$, with its image under $f_{\vec{\epsilon}}$. Let $c_{i}$ for $i=1,...,4$ be the new vertices caused by the cut at the focus-focus critical values $(A,B,C,D)$, respectively. In order to obtain the representative of the polytope invariant we need to compute the slopes of the edges connecting these new introduced vertices with the vertices coming from the elliptic-elliptic values. Let $l_{1}$ be the edge connecting $e_{2}$ to $c_{1}$, $l_{2}$ connect $c_{1}$ to $c_{3}$, $l_{3}$ connect $c_{3}$ to $e_{4}$, $l_{4}$ connect $e_{4}$ to $e_{3}$, $l_{5}$ connect $c_{4}$ to $e_{3}$, $l_{6}$ connect $c_{2}$ to $c_{4}$, $l_{7}$ connect $e_{1}$ to $c_{2}$ and $l_{8}$ connect $e_{1}$ to $e_{2}$. Notice that $l_{8}$ and $l_{4}$ are vertical.
First notice that due to use of an integral affine transformation that leaves a vertical line invariant we may assume that $e_{1}=(0,1)$ and that $l_{7}$ has slope $-1$. Let sl denote the slope of an edge. Using Theorem A.2, we obtain
@@
[…]
Sonja Hohloch &\&& Pedro Santos
University of Antwerp
Department of Mathematics
Middelheimlaan 1
B-2020 Antwerpen, Belgium
E-mail: sonja DOT hohloch AT uantwerpen DOT be
E-mail: pedro DOT santos AT uantwerpen DOT be

### v2

@@ 4.3. A semitoric system induced by the octagon
In De Meulenaere & Hohloch [DH21, Section $3.1$] it is shown that $M=\tilde{L}^{-1}(0)/N$, where
[…]
@@
Points on $(M,\omega)$ are written as equivalence classes of the form $z]=[z_{1},...,z_{8}]$ with $z_{k}=x_{k}+iy_{k}\in\mathbb{C}$ for $1\leq k\leq 8$. The momentum map of the toric system $(M,\omega)$ is given as follows:
@@
(De Meulenaere &\& Hohloch, [DH21, Theorem 1.1]) Let $F=(J,H):M\rightarrow\mathbb{R}^{2}$ be given by
@@
(De Meulenaere &\& Hohloch, [DH21, Theorem 1.2]) Set $F_{t}:=(J,H_{t}):=(J,(1-2t)H+t\gamma X):(M,\omega)\rightarrow\mathbb{R}^{2}$. Then, $(M,\omega,F_{t})_{0\leq t\leq 1}$ is toric for $t=0$, of toric type for $0<t<t^{-},$ semitoric for $t^{-}<t<t^{+}$, and again of toric type for $t^{+}<t\leq 1$ where
$0<t^{-}:=\frac{1}{2(1+24\gamma)}<\frac{1}{2}<t^{+}:=\frac{1}{2(1-24\gamma)}<1.$
@@
(De Meulenaere &\& Hohloch, [DH21, Proposition 3.7]) The focus-focus points of the semitoric systems $(M,\omega,F_{t})$, for $t^{-}<t<t^{+}$ are
@@
(De Meulenaere &\& Hohloch, [DH21, Proposition 1.3]) At $t=\frac{1}{2}$, the system $F_{\frac{1}{2}}$ has precisely two focus-focus fibers, each of which contains precisely two focus-focus points so that each of these two fibers has the shape of a double pinched torus.
@@
of $(\mathbb{C}^{8},\omega_{0})$. Recall that $(M,\omega)$ is obtained as symplectic reduction of $(\mathbb{C}^{8},\omega_{0})$ on a level set $\tilde{L}^{-1}(0)$, see Equation (4.5), by a torus action of $\mathbb{T}^{6}$, see Equation (4.6). Note that $\Psi_{1}$ and $\Psi_{2}$ preserve $\tilde{L}^{-1}(0)$ and are invariant under the action, hence they define symplectomorphisms on $(M,\omega)$. The result follows from considering the projections of $\Psi_{1},\Psi_{2}$$\Psi_{1}\circ\Psi_{2}$ and $\Psi_{2}\circ\Psi_{1}$ to $(M,\omega)$. The symplectomorphisms do not depend on $t$. ∎
@@ 4.3.1. Symplectic topology of the fibers of the toric system
Application of Lemma 2.22: depending on the value one can use the probe with direction $(\pm 1,0)$ or $(0,\pm 1)$, see Figure 4.8. ∎
@@
We present the argument for the fiber over $(\frac{3}{2},\frac{3}{2})$, the other ones are analogous. See Borman [Bor13] and Abreu & Macarini [AM13] for more details on the general method.
  • Notice that $(M,\omega)$ can be viewed as the symplectic reduction of $\mathbb{CP}^{2}\times\mathbb{C}\mathbb{P}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}$ for an appropriate choice of subtorus with associated Lie algebra vectors and appropriate chosen symplectic volume of $\mathbb{CP}^{2}$ and $\mathbb{CP}^{1}$, see Figure 4.9. For more details see Section 2.10 and Abreu & Macarini [AM13].
  • Notice that the central fiber in $(\mathbb{CP}^{1},\omega_{0})$ and the fiber over $(\frac{1}{3},\frac{1}{3})$ in $(\mathbb{CP}^{2},\omega_{FS})$ are stems, and in particular superheavy for any quasi-state, see Polterovich & Rosen [PR14, Proposition $6.1.13$].
  • For each $\mathbb{CP}^{1}$ and for $\mathbb{CP}^{2}$ we choose quasi-states as in Borman [Bor13, Theorem $3.1$] to create a product quasi-state in $\mathbb{CP}^{2}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}$, see Borman [Bor13, Corollary $3.2$], such that the product of the above mentioned fibers is superheavy with respect to this quasi state. Furthermore, this quasi-state satisfies the PB-inequality.
@@
Figure 4.9. $(M,\omega)$ as the symplectic reduction of $\mathbb{CP}^{2}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}$. The dotted lines come in pairs, and each pair corresponds to a reduction by $\mathbb{CP}^{1}$. The triangle corresponds to $\mathbb{CP}^{2}$.
@@ 4.3.2. Symplectic topology of the fibers of the semitoric systems
Figure 4.10. Representative of the polytope invariant for the semitoric system when $\vec{\epsilon}=(-1,1,-1,1)$. The dotted lines represent the eigenrays associated with the focus-focus fibers. The red circles represent the 44 different focus-focus values.
@@ 4.3.3. Case $0<h<1$
Consider the semitoric system given by $(M,\omega,F_{t})$ and $t^{-}<t<t^{+}$ with height invariant $0<h<1$. Then there exist $5$ nondisplaceable fibers. These are the fibers over the points $\{(1,1),(1,2),(\frac{3}{2},\frac{3}{2}),(2,1),(2,2)\}$ in the polytope invariant given by Figure 4.10. All other fibers are displaceable. In particular the focus-focus fibers are displaceable.
@@ 4.3.5. The case $h=1$
  • Applying Lemma 2.22 to Figure 4.16 2.22 we conclude that every fiber of the form $(h_{\vec{\epsilon}}\circ F_{t})^{-1}(1,b)$ with $1<b<2$, where $h_{\vec{\epsilon}}$ is the straightening homeomorphism associated with the polytope of Figure 4.16, is displaceable.
@@ Appendix A Polytope invariant for the semitoric system induced by the octagon
Let $(M,\omega)$ be a closed symplectic manifold of dimension $2n$. The Liouville measure on $(M,\omega)$ is the measure induced by the volume form $\frac{1}{(2\pi)^{n}}\frac{\omega^{n}}{n!}$. The Duistermaat-Heckman measure $\mu_{J}$ for a Hamiltonian $J:M\rightarrow\mathbb{R}$ inducing an $\mathbb{S}^{1}$-action is defined as $\mu_{J}([a,b])=\text{vol}(J^{-1}([a,b]))$ where vol is taken with respect to the Liouville measure in $M$. By Duistermaat & Heckman [DH82]
@@
where the density $\rho_{J}(x)$, sometimes called the Duistermaat-Heckman function, is a continuous function, equal to the symplectic volume of the reduced orbifold $J^{-1}(x)/\mathbb{S}^{1}$.
Let $J$ be the momentum map of an effective Hamiltonian $\mathbb{S}^{1}$-action on $(M,\omega)$. For each subgroup $G\subset\mathbb{S}^{1}$, let $M^{G}$ be the set of points in $M$ whose stabilizer is $G$. The connected components of $M^{\mathbb{S}^{1}}$ are symplectic submanifolds, hence either points or surfaces, see Karshon [Kar99] for more details.
@@
(Chaperon, [Cha83]) Let $(M,\omega,J)$ be as above. For each $p\in M^{\mathbb{S}^{1}}$ there exist neighborhoods $U\subset M$ of $p$, $U_{0}\subset\mathbb{C}^{2}$ of $(0,0)$, and a symplectomorphism $\Psi:(U,\omega)\rightarrow(U_{0},\omega_{0})$, where $\omega_{0}=\frac{i}{2}(dz_{1}\wedge d\overline{z}_{1}+dz_{2}\wedge d\overline{z}_{2})$ making the following diagram commute
@@
(Vũ Ngǫc, [Vu07, Theorem 3.8]) Let $(M,\omega,\Phi)$ be a semitoric system and $f_{\vec{\epsilon}}$ a straightening homeomorphism associated with a representative of the polytope invariant. Furthermore, let $x\in\mathbb{R}$ be a regular value of $J$. If $\alpha^{+}(x)$ (resp. $\alpha^{-}(x)$) denotes the slope of the top (resp. bottom) boundary of the polytope $f_{\vec{\epsilon}}\circ\Phi(M)$, then the derivative of the Duistermaat-Heckman function is
@@
where $a^{\pm},b^{\pm}$ are the isotropy weights for the $\mathbb{S}^{1}$-action at the corresponding vertices.
With Theorem A.2, we are able to compute representatives of the polytope invariant (see Section 2.8) for the semitoric system given in Section 4.3. We focus on the case of $4$ focus-focus values $(A,B,C,D)$. Furthermore, recall that there are $4$ elliptic-elliptic values. We label them $e_{1}$, $e_{2}$, $e_{3}$, $e_{4}$. We now compute a representative of the polytope invariant for the case $\vec{\epsilon}=(-1,1,-1,1)$, the other cases are analogous.
@@
Consider the image of $F(M)$ under the straightening homeomorphism $f_{\vec{\epsilon}}$. By abuse of notation we identify each $e_{i}$, $i=1,..,4$, with its image under $f_{\vec{\epsilon}}$. Let $c_{i}$ for i=1,,4i=1,...,4 be the new vertices caused by the cut at the focus-focus critical values $(A,B,C,D)$, respectively. In order to obtain the representative of the polytope invariant we need to compute the slopes of the edges connecting these new introduced vertices with the vertices coming from the elliptic-elliptic values. Let $l_{1}$ be the edge connecting $e_{2}$ to $c_{1}$, $l_{2}$ connect $c_{1}$ to $c_{3}$, $l_{3}$ connect $c_{3}$ to $e_{4}$, $l_{4}$ connect $e_{4}$ to $e_{3}$, $l_{5}$ connect $c_{4}$ to $e_{3}$, $l_{6}$ connect $c_{2}$ to $c_{4}$, $l_{7}$ connect $e_{1}$ to $c_{2}$ and $l_{8}$ connect $e_{1}$ to $e_{2}$. Notice that $l_{8}$ and $l_{4}$ are vertical.
First notice that due to use of an integral affine transformation that leaves a vertical line invariant we may assume that $e_{1}=(0,1)$ and that $l_{7}$ has slope 1-1. Let sl denote the slope of an edge. Using Theorem A.2, we obtain